OpenStudy (anonymous):

Find 6 distinct points on the circle with center (4,1) and circumference 3. Show work.

OpenStudy (anonymous):

circumference is 3??

OpenStudy (anonymous):

guess you need the radius, \[C=2\pi r\] \[3=2\pi r\] \[r=\frac{2\pi}{3}\] and equation for the circle is therefore \[(x-4)^2+(y-2)^2=\frac{4\pi^2}{9}\]

OpenStudy (anonymous):

typo there, equation is \[(x-4)^2+(y-1)^2=\frac{4\pi^2}{9}\]

OpenStudy (anonymous):

pick an x, find the y lather, rinse, repeat